OpenStudy (pokemon23):

Suppose that you are selling sweatshirts for a class fund-raiser.The wholesaler charges you $8 for each sweatshirt. You charge $16 for each sweatshirt. Find the percent of increase.

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

Percent increase = (new value - original value)/original value * 100 In this case, it will be (16 - 8) / 8 * 100 = 8/8 * 100 = 100%